OpenStudy (anonymous):

Volumes of Revolution Help. Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis. y = (1/x^2), y=0, x=3, x=5 about y=-1 I'm not sure what to do with this problem. I graphed it and it became more confusing once i saw it.

OpenStudy (anonymous):

You are bounded by x=3 and x=5. Are you using the shell method, the disk method, or cross sections?

OpenStudy (anonymous):

I would recommend the disk/washer method to find the volume. Draw a rectangle perpendicular to y=-1 and touching y=1/x*2.

OpenStudy (anonymous):

Im not sure which method to use

OpenStudy (anonymous):

Then your integral becomes:\[\int\limits_{3}^{5}\ [(\frac{ 1 }{x ^{2} }+1)^{2}-1^{2}]dx\]

OpenStudy (anonymous):

Why is 1 added to (1/x^2)?

OpenStudy (anonymous):

Because when you are revolving about the line y=a, the resulting radius is f(x)-a. In this case, a=-1, so the full radius from the line y=-1 to the function is f(x)-(1)

OpenStudy (anonymous):

oops, that should be f(x)-(-1))

OpenStudy (anonymous):

ah okay i see that

OpenStudy (anonymous):

Now you get a much easier integral when the 1's cancel after squaring out the first part.

OpenStudy (anonymous):

So, your integral becomes:\[\int\limits_{3}^{5}\left( \frac{ 1 }{ x ^{4} } +\frac{ 2 }{ x ^{2} }\right)dx\]

OpenStudy (anonymous):

Now you can integrate and get: \[-\frac{ 1 }{ 3 } x ^{-3}-2x ^{-1}\] Evaluating at the given values, you have: \[-\frac{ 1 }{ 3}(5)^{-3}-2(5)^{-1}-[-\frac{ 1 }{ 3 }3^{-3}-2(3)^{-1}\]

OpenStudy (anonymous):

Finally, you can simplify that mess and get \[-\frac{ 1 }{375 }-\frac{ 2 }{ 5 }+\frac{ 1 }{ 81 }+\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

I think you can see from here that you have to get an LCD to finish. Sorry, but not used to the equation editor so taking me awhile to type out replies.

OpenStudy (anonymous):

By the way, I forgot the pi at the beginning, so don't forget to multiply your final answer by pi.

OpenStudy (anonymous):

Thank you so much. You've been a huge help!

OpenStudy (anonymous):

No prb.